OpenStudy (sasogeek):

Two smooth spheres \(\large A \) and \(\large B \) collide. The mass of \(\large A \) is four times that of \(\large B. \) The final velocities of \(\large A \) and \(\large B \) after collision are \(\large i +2j \) and \(\large -4i+7j \) respectively. If the initial velocity of \(\large A \) is \(\large -2i+5j \) and the linear momentum is conserved, find the initial velocity of \(\large B. \)

OpenStudy (bahrom7893):

this belongs in physics, though im sure that if you post all formulas we'll figure this out.. i think it was something like, u calculate momentum which was something like mass times something else, and then u calculate final momentum, equate the equations, because momentum is conserved, and you're done

OpenStudy (bahrom7893):

Sorry i haven't looked at physics in two years.

OpenStudy (sasogeek):

well momentum conserved after collision is \(\large M_{1}U_{1}+M_{2}U_{2}=M_{1}V_{1}+M_{2}V_{2} \) where M=mass , U=initial velocity and V=final velocity

OpenStudy (bahrom7893):

oh okay this is simple then

OpenStudy (bahrom7893):

Ma = 4Mb V_fa = i+2j V_fb = -4i+7j V_ia = -2i+5j V_ib - ?

OpenStudy (bahrom7893):

Now set up the equation: Ma*V_ia + Mb*V_ib = Ma*V_fa + Mb*V_fb

OpenStudy (bahrom7893):

4Mb*(-2i+5j)+Mb*V_ib = 4Mb*(i+2j)+Mb*(-4i+7j)

OpenStudy (bahrom7893):

While i compute this further, double check that i made all correct substitutions.

OpenStudy (bahrom7893):

Mb*V_ib = 4Mb*(i+2j)+Mb*(-4i+7j) - 4Mb*(-2i+5j) Divide everything by Mb V_ib = 4*(i+2j)+(-4i+7j)-4*(-2i+5j) V_ib = 4i + 8j - 4i + 7j +8i - 20j V_ib = 8j + 7j + 8i - 20j V_ib = 15j - 20j + 8i V_ib = 8i - 5j

OpenStudy (sasogeek):

thanks :) i had to go on an errand :P just got back ;)